OpenStudy (he66666):

A plane passes through the points P(-2, 3, 1), Q(-2, 3, 2), and R(1, 0, 1). Using vector QR and one other direction vector, write a second vector equation for this plane. -> Can I choose vector QP as my other director vector and have r=(-2, 3, 2) +s(3, -3, -1) +t(0, 0, -1) as my answer? Cause the answer says it's r=(-2, 3, 2) +s(3, -3, -1) +t(0, 0, 1)..

OpenStudy (amistre64):

vector equation for the plane eh.....

OpenStudy (amistre64):

the vector equation of a plane that i am used to consists of a normal attached to a point.....

OpenStudy (amistre64):

the vector equation of a surface is made up of parametric equation right?

OpenStudy (he66666):

Yes, but the textbook didn't solve it using the normal- just by the three points :S

OpenStudy (amistre64):

when we add vectors; i spose that is what it is refering to...maybe?

OpenStudy (amistre64):

P(x1,y1,z1); v<a,b,c> x = x1- at y = y1- bt z = z1- ct f(x,y,z) = (x1-at)i + (y1-bt)j + (z1-ct)k then?

OpenStudy (amistre64):

well, thats subtracting vector if anthing :)

OpenStudy (he66666):

oh okay, I think I get it. Thanks amistre :D

OpenStudy (amistre64):

youre welcome :)